Circular No. 6331

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

4U 1608–52 F. E. Marshall, Laboratory for High Energy Astrophysics, Goddard

Space Flight Center (LHEA/GSFC); L. Angelini, Universities Space Research Association and LHEA/GSFC; and the X-ray Timing Explorer (XTE) Science Operations Center Team report the detection of the recurrent soft-x-ray transient 4U 1608–52 (cf. IAUC 2859, 3090) in outburst. 4U 1608–52 is a low-mass x-ray binary that also exhibits type-I x-ray bursts. During a brief scan of the Galactic plane with the Proportional Counter Array (PCA) onboard XTE, 4U 1608–52 was seen on Feb. 23.883 UT with a countrate of ∼ 5700 counts/s (∼ 44 percent of the rate for the Crab). The 3- to 20-keV spectrum of the source can be qualitatively described by a thermal bremsstrahlung model with kT = 5 keV and absorption with column density N_{H} = 3 × 22 atoms cm^{−2}. About ten outbursts from this source have been observed since 1970 (Lochner and Roussel-Dupré 1994, Ap.J. 435, 840). Flaring from the proposed optical counterpart (QX Nor; V = 18.2) was seen during the 1977 outburst (Grindlay and Liller 1978, Ap.J. 220, L127). This neutron-star transient is of particular interest because it sometimes shows some characteristics similar to the blackhole candidate Cyg X-1. Observations at other wavelengths and by other x-ray observatories are encouraged.”

NOVA CASSIOPEIAE 1995 Visual magnitude estimates (cf. IAUC 6308), provided in part by J.

Mattei, AAVSO: Feb. 5.11 UT, 8.9 (P. Collins, Scottsdale, AZ); 12.08, 9.2 (A. Dill, Wichita, KS); 14.87, 8.5 (M. V. Zanotta, Laino, Italy); 16.82, 8.4 (A. Pereira, Cabo da Roca, Portugal); 18.12, 8.1 (W. G. Dillon, Missouri City, TX); 25.17, 8.3 (C. E. Spratt, Victoria, BC); 26.76, 8.5 (P. Schmeer, Bischmisheim, Germany).

COMET C/1996 B1 (SZCZEPANSKI) Total visual magnitude and coma-diameter estimates (cf. IAUC 6306):

Feb. 10.16 UT, 8.7, 7′.5 (J. E. Bortle, Stormville, NY, 20×80 binoculars); 17.18, 8.2, 12′ (R. Haver, Frasso Sabino, Italy, 15×80 binoculars); 21.32, 8.1, 12′ (G. Kronk, Troy, IL, 20×80 binoculars); 24.31, 7.4, 30′ (C. S. Morris, Lockwood Valley, 10×50 binoculars); 25.88, 7.2, ∼ 20′ (K. Hornoch, Lelekovice, Czech Republic, 10×80 binoculars); 29.38, 8.0, 14′ (Kronk).

1996 March 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green